# Convergence and long-time behavior of finite volumes for a generalized Poisson-Nernst-Planck system with cross-diffusion and size exclusion

### v1

@@ 3.3. Further uniform estimates on the discrete solution
$\sum_{i=1}^{I}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}\frac{d_{\sigma}}{m_{\sigma}}\left|F_{i,K\sigma}^{\text{diff},n}\right|^{2}$$\leq$ $C\sum_{i=1}^{I}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left|\sqrt{u_{i,K}^{n}u_{0,K}^{n}}-\sqrt{u_{i,L}^{n}u_{0,L}^{n}}\right|^{2}$
$+C\sum_{i=1}^{I}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}|u_{0,L}^{n}-u_{0,K}^{n}|^{2},$
thanks to the uniform bounds on $(U^{n}_{K})_{K\in\mathcal{T},n\leq 1}$. Lemma 3.7 provides the desired $L^{2}_{\text{loc}}(L^{2})$ bound on $F_{i,K\sigma}^{\text{diff},n}$, concluding the proof of Lemma 3.8. ∎
From Lemma 3.8, we deduce the following discrete $L^{2}_{\text{loc}}((H^{1})^{\prime})$ estimate:
@@
There exists $C$ depending only on $\Omega$, $\phi^{D}$, $\lambda$, $f$, ${(z_{i})}_{i}$, ${(D_{i})}_{i}$ and $\zeta_{\mathcal{T}}$ such that, for all $i=0,\dots,I$ and all $\varphi_{\mathcal{T},\tau}=\sum_{K\in\mathcal{T}}\sum_{n=1}^{N}\varphi_{K}^{n}\mathbbm{1}_{(t^{n-1},t^{n}]\times K}$, one has
(55) $\sum_{n=1}^{N}\sum_{K\in\mathcal{T}}(u_{i,K}^{n}-u_{i,K}^{n-1})\varphi_{K}^{n}\leq C\left(1+t^{N}\right)^{1/2}\left(\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left(\varphi_{K}^{n}-\varphi_{K\sigma}^{n}\right)^{2}\right)^{1/2}.$
@@
Let us first establish (55) for $i\geq 1$. Multiplying (25) by $\tau^{n}\varphi_{K}^{n}$, summing over $K\in\mathcal{T}$ and $n\in\{1,\dots,N\}$, and performing a discrete integration by parts on the contribution of the fluxes provides
(56) $\sum_{n=1}^{N}\sum_{K\in\mathcal{T}}m_{K}(u_{i,K}^{n}-u_{i,K}^{n-1})\varphi_{K}^{n}=\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}F_{i,K\sigma}^{n}(\varphi_{K\sigma}^{n}-\varphi_{K}^{n}).$
Applying Cauchy-Schwarz inequality to the right-hand side then using Lemma 3.8 provides the desired result. The recovery of the estimate for $i=0$ then directly follows from the definition (27) of $u_{0,K}^{n}$ and from (55) for $i\geq 1$. ∎
@@ 4. Convergence of the schemes
As usual in the analysis of finite volume schemes, we also need to handle quantities attached to the faces $\sigma\in\mathcal{E}$. To this end, we introduce the so-called diamond mesh of $\Omega$ by associating a diamond cell $\omega_{\sigma}$ to all $\sigma\in\mathcal{E}$. More precisely, $\omega_{\sigma}$ is the convex hull of $\{x_{K},x_{L},\sigma\}$ if $\sigma=K|L\in\mathcal{E}_{\text{int}}$, and $\omega_{\sigma}=\operatorname{conv}\{x_{K},\sigma\}$ if $\sigma\in\mathcal{E}_{\text{ext}}\cap\mathcal{E}_{K}$, see Figure 1 for an illustration. The Lebesgue measure $m_{\omega_{\sigma}}$ of $\omega_{\sigma}$ is then given by
@@
$F_{i,\mathcal{E}_{\ell},\tau_{\ell}}(t,x)\underset{\ell\to+\infty}{\longrightarrow}F_{i}\quad\text{weakly in}\;L^{2}_{\text{loc}}(\mathbb{R}_{+}\times\overline{\Omega})^{d}.$
@@ 4.1. Compactness of the approximate solutions
There exist functions $u_{0}\in L^{\infty}(\mathbb{R}_{+}\times\Omega)$ with $\sqrt{u_{0}}\in L^{2}_{\text{loc}}(\mathbb{R}_{+};H^{1}(\Omega))$, $u_{1},\dots,u_{n}\in L^{\infty}(\mathbb{R}_{+}\times\Omega)$ with $u_{i}\sqrt{u_{0}}\in L^{2}_{\text{loc}}(\mathbb{R}_{+};H^{1}(\Omega))$, and $\phi\in L^{\infty}(\mathbb{R}_{+}\times\Omega)$ with $\phi-\phi^{D}\in L^{\infty}(\mathbb{R}_{+};V)$ such that, up to a subsequence, as $\ell\to+\infty$, there holds
(60) $u_{i,\mathcal{T},\tau}\rightarrow u_{i}$ in the L(+×Ω)-weak- sense, $\quad\text{in the $L^{\infty}(\mathbb{R}_{+}\times\Omega)$-weak-$\star$ sense},\;1\leq i\leq I,$
(61) $u_{0,\mathcal{T},\tau}\rightarrow u_{0}$ in the L(+×Ω)-weak- sense and a.e. in +×$\quad\text{in the $L^{\infty}(\mathbb{R}_{+}\times\Omega)$-weak-$\star$ sense and a.e. in $\mathbb{R}_{+}\times\Omega$},$
(62) $\phi_{\mathcal{T},\tau}\rightarrow\phi$ in the L(+×Ω)-weak- sense and a.e. in +×$\quad\text{in the $L^{\infty}(\mathbb{R}_{+}\times\Omega)$-weak-$\star$ sense and a.e. in $\mathbb{R}_{+}\times\Omega$}.$
@@
[…]
@@
(65) $\nabla_{\mathcal{T},\tau}\sqrt{u_{0,\mathcal{T},\tau}}\rightharpoonup\nabla\sqrt{u_{0}}$ weakly in $\quad\text{weakly in }L^{2}_{\text{loc}}(\mathbb{R}_{+};L^{2}(\Omega)^{d}),$
(66) $\nabla_{\mathcal{T},\tau}[u_{0,\mathcal{T},\tau}u_{i,\mathcal{T},\tau}]\rightharpoonup\nabla[u_{0}u_{i}]$ weakly in $\quad\text{weakly in }L^{2}_{\text{loc}}(\mathbb{R}_{+};L^{2}(\Omega)),$
(67) $\nabla_{\mathcal{T},\tau}\phi_{\mathcal{T},\tau}\rightharpoonup\nabla\phi$ weakly in $\quad\text{weakly in }L^{2}_{\text{loc}}(\mathbb{R}_{+};L^{2}(\Omega)),$
@@
The $L^{\infty}(\mathbb{R}_{+}\times\Omega)$-weak-$\star$ convergences (60)–(62) are the consequences of the uniform bounds on $u_{i,\mathcal{T},\tau}$ and $\phi_{\mathcal{T},\tau}$ in $L^{\infty}(\mathbb{R}_{+}\times\Omega)$, thanks to the Banach-Alaoglu Theorem. The almost everywhere convergence (61) follows from strong convergence given by the Aubin-Lions lemma, see [26] for a general presentation of the lemma, and [7, Lemma 9] for its application in our context. It can be applied thanks to the uniform estimates of Lemma 3.7 and Corollary 3.9. These estimates also yields the weak convergence in $L^{2}_{\text{loc}}(\mathbb{R}_{+};L^{2}(\Omega))$ of $\nabla_{\mathcal{T},\tau}u_{0,\mathcal{T},\tau}$ towards $\nabla u_{0}$. Concerning the point-wise convergence of $\phi_{\mathcal{T},\tau}$ towards $\phi$, it can be proven by using the discrete $L^{2}_{\text{loc}}((H^{1})^{\prime})$ estimate on time increments of the right-hand side of the discrete Poisson equation (24) that follows from Corollary 3.9. As the proof is fully similar to the one of [14, Proposition 4.2], we do not provide details here. The weak compactness property (67) is also established in [14, Proposition 4.2].
@@
In order to prove (65), one can either make use of the nonlinear Aubin Simon theorem of [1], or directly remark that, thanks to Lemma 3.7, the vector field $\nabla_{\mathcal{T}}\sqrt{u_{0,\mathcal{T},\tau}}$ is uniformly bounded in $L^{2}_{\text{loc}}(\mathbb{R}_{+};L^{2}(\Omega)^{d})$. We deduce from this boundedness that there exists $G_{0}\in L^{2}_{\text{loc}}(\mathbb{R}_{+};L^{2}(\Omega)^{d})$ such that, up to a subsequence, $\nabla_{\mathcal{T}}\sqrt{u_{0,\mathcal{T},\tau}}$ tends to $G_{0}$ weakly. As $\sqrt{u_{0,\mathcal{T},\tau}}$ converges point-wise (and thus in $L^{1}_{\text{loc}}(\mathbb{R}_{+}\times\Omega)$) towards $\sqrt{u_{0}}$ owing to (61), the weak consistency of the inflated gradient $\nabla_{\mathcal{T}}$, see for instance [17, 20], allows to show that $G_{0}=\nabla\sqrt{u_{0}}$.
@@
Regarding the point-wise convergences (63) and (64), of $u_{i,\mathcal{T},\tau}\sqrt{u_{0,\mathcal{T},\tau}}$, respectively, we can conclude by applying the discrete Aubin-Lions lemma of ”degenerate” type ([7, Lemma 10]). To prove (63), we use it with $y_{\ell}=\sqrt{u_{0,\mathcal{T}_{\ell},\tau_{\ell}}}$ and $z_{\ell}=u_{i,\mathcal{T}_{\ell},\tau_{\ell}}$, making advantages of the uniform estimate (55), (47) and (48). Whereas, with $y_{\ell}=u_{0,\mathcal{T}_{\ell},\tau_{\ell}}$ and $z_{\ell}=u_{i,\mathcal{T}_{\ell},\tau_{\ell}}$, we conclude (64) similarly, replacing (48) by (49).
@@
It finally remains to prove (66) for $i=1,\dots,I$. As a consequence of (49), $\nabla_{\mathcal{T}}[u_{i,\mathcal{T},\tau}u_{0,\mathcal{T},\tau}]$ is uniformly bounded in $L^{2}_{\text{loc}}(\mathbb{R}_{+}\times\overline{\Omega})^{d}$. Hence there exists $G_{i}$ such that, up to a subsequence,
@@
$\nabla_{\mathcal{T}}[u_{i,\mathcal{T},\tau}u_{0,\mathcal{T},\tau}]\rightharpoonup G_{i}\quad\text{weakly in }L^{2}_{\text{loc}}(\mathbb{R}_{+}\times\overline{\Omega})^{d}.$

### v2

@@ 3.3. Further uniform estimates on the discrete solution
$\sum_{i=1}^{I}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}\frac{d_{\sigma}}{m_{\sigma}}\left|F_{i,K\sigma}^{\text{diff},n}\right|^{2}$ $\leq$ $C\sum_{i=1}^{I}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left|\sqrt{u_{i,K}^{n}u_{0,K}^{n}}-\sqrt{u_{i,L}^{n}u_{0,L}^{n}}\right|^{2}$
$+C\sum_{i=1}^{I}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}|u_{0,L}^{n}-u_{0,K}^{n}|^{2},$
thanks to the uniform bounds on $(U^{n}_{K})_{K\in\mathcal{T},n\leq 1}$. Lemma 3.8 provides the desired $L^{2}_{\rm loc}(L^{2})$ bound on $F_{i,K\sigma}^{\text{diff},n}$, concluding the proof of Lemma 3.9. ∎
From Lemma 3.9, we deduce the following discrete $L^{2}_{\rm loc}((H^{1})^{\prime})$ estimate:
@@
There exists $C$ depending only on $\Omega$, $\phi^{D}$, $\lambda$, $f$, ${(z_{i})}_{i}$, ${(D_{i})}_{i}$, and $\zeta_{\mathcal{T}}$ such that, for all $i=0,\dots,I$ and all $\varphi_{\mathcal{T},\tau}=\sum_{K\in\mathcal{T}}\sum_{n=1}^{N}\varphi_{K}^{n}\mathbbm{1}_{(t^{n-1},t^{n}]\times K}$, one has
(55) $\sum_{n=1}^{N}\sum_{K\in\mathcal{T}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m_{K}}(u_{i,K}^{n}-u_{i,K}^{n-1})\varphi_{K}^{n}\leq C\left(1+t^{N}\right)^{1/2}\left(\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(\varphi_{K}^{n}-\varphi_{K\sigma}^{n}\right)^{2}\right)^{1/2}.$
@@
Let us first establish (55) for $i\geq 1$. Multiplying (25) by $\tau^{n}\varphi_{K}^{n}$, summing over $K\in\mathcal{T}$ and $n\in\{1,\dots,N\}$, and performing a discrete integration by parts on the contribution of the fluxes, namely $\sum_{K\in\mathcal{T}}\sum_{\sigma\in\mathcal{E}_{K}}F_{i,K\sigma}^{n}\varphi_{K}^{n}=\sum_{\sigma\in\mathcal{E}_{\rm int}}F_{i,K\sigma}^{n}(\varphi_{K\sigma}^{n}-\varphi_{K}^{n})$ for all $i=1,\dots,I$, provides
(56) $\sum_{n=1}^{N}\sum_{K\in\mathcal{T}}m_{K}(u_{i,K}^{n}-u_{i,K}^{n-1})\varphi_{K}^{n}=\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}F_{i,K\sigma}^{n}(\varphi_{K\sigma}^{n}-\varphi_{K}^{n}).$
Applying Cauchy–Schwarz inequality to the right-hand side then using Lemma 3.9 provides the desired result. The recovery of the estimate for $i=0$ then directly follows from the definition (27) of $u_{0,K}^{n}$ and from (55) for $i\geq 1$. ∎
@@ 4. Convergence of the schemes
As usual in the analysis of finite volume schemes, we also need to handle quantities attached to the faces $\sigma\in\mathcal{E}$. To this end, we introduce the so-called diamond mesh of $\Omega$ by associating a diamond cell $\omega_{\sigma}$ to every $\sigma\in\mathcal{E}$. More precisely, $\omega_{\sigma}$ is the convex hull of $\{x_{K},x_{L},\sigma\}$ if $\sigma=K|L\in\mathcal{E}_{\rm int}$, and $\omega_{\sigma}=\operatorname{conv}\{x_{K},\sigma\}$ if $\sigma\in\mathcal{E}_{\text{ext}}\cap\mathcal{E}_{K}$, see Figure 1 for an illustration. The Lebesgue measure $m_{\omega_{\sigma}}$ of $\omega_{\sigma}$ is then given by
@@
$F_{i,\mathcal{E}_{\ell},\tau_{\ell}}(t,x)\underset{\ell\to+\infty}{\longrightarrow}F_{i}\quad\text{weakly in}\;L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\overline{\Omega})^{d}.$
@@ 4.1. Compactness of the approximate solutions
There exist functions $u_{0}\in L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)$ with $\sqrt{u_{0}}\in L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};H^{1}(\Omega))$, $u_{1},\dots,u_{n}\in L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)$ with $u_{i}\sqrt{u_{0}}\in L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};H^{1}(\Omega))$, and $\phi\in L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)$ with $\phi-\phi^{D}\in L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};V)$ such that, up to a subsequence, as $\ell\to+\infty$, there holds
(60) $u_{i,\mathcal{T},\tau}\rightarrow u_{i}$ $\quad\text{in the $L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)$-weak-$\star$ sense},\;1\leq i\leq I,$
(61) $u_{0,\mathcal{T},\tau}\rightarrow u_{0}$ $\quad\text{in the $L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)$-weak-$\star$ sense and a.e. in ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega$},$
(62) $\phi_{\mathcal{T},\tau}\rightarrow\phi$ $\quad\text{in the $L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)$-weak-$\star$ sense and a.e. in ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega$}.$
@@
(63) $u_{i,\mathcal{T},\tau}\sqrt{u_{0,\mathcal{T},\tau}}\rightarrow u_{i}\sqrt{u_{0}}$ $\quad\text{a.e. in }{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega,$
(64) $u_{0,\mathcal{T},\tau}u_{i,\mathcal{T},\tau}\rightarrow u_{0}u_{i}$ $\quad\text{a.e. in }{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega,$
@@
(65) $\nabla_{\mathcal{T},\tau}\sqrt{u_{0,\mathcal{T},\tau}}\rightharpoonup\nabla\sqrt{u_{0}}$ $\quad\text{weakly in }L^{2}_{{\rm loc}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};L^{2}(\Omega)^{d}),$
(66) $\nabla_{\mathcal{T},\tau}[u_{0,\mathcal{T},\tau}u_{i,\mathcal{T},\tau}]\rightharpoonup\nabla[u_{0}u_{i}]$ $\quad\text{weakly in }L^{2}_{{\rm loc}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};L^{2}(\Omega)),$
(67) $\nabla_{\mathcal{T},\tau}\phi_{\mathcal{T},\tau}\rightharpoonup\nabla\phi$ $\quad\text{weakly in }L^{2}_{{\rm loc}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};L^{2}(\Omega)),$
@@
The $L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)$-weak-$\star$ convergences (60)–(62) are the consequences of the uniform bounds on $u_{i,\mathcal{T},\tau}$ and $\phi_{\mathcal{T},\tau}$ in $L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)$, thanks to the Banach–Alaoglu Theorem. The almost everywhere convergence (61) follows from strong convergence given by the Aubin–Lions lemma, see [29] for a general presentation of the lemma, and [7, Lemma 9] for its application in our context. It can be applied thanks to the uniform estimates of Lemma 3.8 and Corollary 3.10. These estimates also yield the weak convergence in $L^{2}_{{\rm loc}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};L^{2}(\Omega))$ of $\nabla_{\mathcal{T},\tau}u_{0,\mathcal{T},\tau}$ towards $\nabla u_{0}$. Concerning the point-wise convergence of $\phi_{\mathcal{T},\tau}$ towards $\phi$, it can be proven by using the discrete $L^{2}_{\rm loc}((H^{1})^{\prime})$ estimate on time increments of the right-hand side of the discrete Poisson equation (24) that follows from Corollary 3.10. As the proof is fully similar to the one of [14, Proposition 4.2], we do not provide details here. The weak compactness property (67) is also established in [14, Proposition 4.2].
@@
In order to prove (65), one can either make use of the non-linear Aubin–Simon theorem of [1], or directly remark that, thanks to Lemma 3.8, the vector field $\nabla_{\mathcal{T}}\sqrt{u_{0,\mathcal{T},\tau}}$ is uniformly bounded in $L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};L^{2}(\Omega)^{d})$. We deduce from this boundedness that there exists $G_{0}\in L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};L^{2}(\Omega)^{d})$ such that, up to a subsequence, $\nabla_{\mathcal{T}}\sqrt{u_{0,\mathcal{T},\tau}}$ tends to $G_{0}$ weakly. As $\sqrt{u_{0,\mathcal{T},\tau}}$ converges point-wise (and thus in $L^{1}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)$) towards $\sqrt{u_{0}}$ owing to (61), the weak consistency of the inflated gradient $\nabla_{\mathcal{T}}$, see for instance [17, 22], allows to show that $G_{0}=\nabla\sqrt{u_{0}}$.
@@
Regarding the point-wise convergences (63) and (64), of $u_{i,\mathcal{T},\tau}\sqrt{u_{0,\mathcal{T},\tau}}$, respectively, we can conclude by applying [7, Lemma 10. (Discrete Aubin–Lions of “degenerate” type)]. To prove (63), we use it with $y_{\ell}=\sqrt{u_{0,\mathcal{T}_{\ell},\tau_{\ell}}}$ and $z_{\ell}=u_{i,\mathcal{T}_{\ell},\tau_{\ell}}$, taking advantage of the uniform estimate (55), (47), and (48). Whereas, with $y_{\ell}=u_{0,\mathcal{T}_{\ell},\tau_{\ell}}$ and $z_{\ell}=u_{i,\mathcal{T}_{\ell},\tau_{\ell}}$, we conclude (64) similarly, replacing (48) by (49).
@@
It finally remains to prove (66) for $i=1,\dots,I$. As a consequence of (49), $\nabla_{\mathcal{T}}[u_{i,\mathcal{T},\tau}u_{0,\mathcal{T},\tau}]$ is uniformly bounded in $L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\overline{\Omega})^{d}$. Hence there exists $G_{i}$ such that, up to a subsequence,
@@
$\nabla_{\mathcal{T}}[u_{i,\mathcal{T},\tau}u_{0,\mathcal{T},\tau}]\rightharpoonup G_{i}\quad\text{weakly in }L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\overline{\Omega})^{d}.$